Comparative data on three companies in the same service industry are given below:

Required:

2. Fill in the missing information. (Round the "Turnover" and "ROI" answers to 2 decimal places.)

Company
A B C
Sales $3,315,000   $2,776,000  
Net operating income $464,100 $388,640
Average operating assets $1,950,000 $3,000,000  
Margin % % 3 %
Turnover 2.00
Return on investment (ROI) % 11.20 % %
Answer #1
Company
A B C

Sales

3315000

2776000

6000000

[ Average operating assets * Turnover = 3000000 * 2.00 ]

Net operating income

464100

388640

180000

[ Sales * Margin = 6000000 * 3% ]

Average operating assets

1950000

3470000

[ Net operating income / ROI = 388640 / 11.20% ]

3000000

Margin

14%

[ Net operating income / Sales = 464100 / 3315000 ]

14%

[ Net operating income / Sales = 388640 / 2776000 ]

3%

Turnover

1.70

[ Sales / Average operating assets = 3315000 / 1950000 ]

0.80

[ Sales / Average operating assets = 2776000 / 3470000 ]

2.00

Return on Investment (ROI)

23.80%

[ Net operating income / Average operating assets = 464100 / 1950000 ]

11.20%

6.00%

[ Net operating income / Average operating assets ]